# White Men Lead, Black Women Help: Uncovering Gender, Racial, and Intersectional Bias in Language Agency

### v1

@@ 5 Conclusion
In this work, we propose and study gender, racial, and intersectional biases in language agency in a wide scope of human-written and LLM-generated texts. First, we contribute the Language Agency Classification (LAC) dataset for training accurate classifiers to measure language agency. We then utilize an agency classifier to reveal critical language agency biases in $6$ human-written and LLM-generated datasets, including biographies, professor reviews, and letters of recommendation. Our findings show that (1) language agency is effective in capturing gender, racial, and intersectional biases in texts, (2) while both are biased, LLM-generated texts demonstrate notably higher levels of language agency bias than human-written texts, and (3) severe biases in language agency target females and Black individuals, especially for Black females. Our findings point towards the imminence to further explore biases in language agency, and forewarn potential social risks of using LLM-generated texts without scrutiny.
Limitations
We identify some limitations of our study. First, due to the limited information within the datasets available for our study, we were only able to consider the binary gender and $4$ racial groups for bias analyses. However, we note that it is important and significant for further works to extend the investigation of the fairness problem in our study to other gender and racial minority groups. Second, our study only includes experiments on investigating language agency biases in personal biographies, professor reviews, and reference letters, among which $4$ were synthesized and only $2$ contained human-written texts. We encourage future studies to extend the exploration of language agency biases in broader domains of human-written texts. Lastly, due to cost and resource constraints, we were not able to further extend our experiments to larger synthesized datasets. Future works should be devoted to comprehensively evaluating biases from various data sources.
Ethics Statement
Experiments in this study incorporate language models that were pre-trained on a wide range of text from the internet and have been shown to learn or amplify biases from the data used. Since we utilize a language model to synthesize a language agency classification dataset, we adopt a number of methods to prevent potential harm and bias propagation: (1) we prompt the model to paraphrase each input into an agentic version and a communal version, ensuring the balance in the preliminary generated dataset, and (2) we invite expert annotators to re-annotate the generated data, to verify and ensure the quality of the final dataset used to train language agency classifiers. We encourage future extensions of our works to also consider this factor in their research, so as to draw reliable and trustworthy research conclusions.
Appendix A Language Agency Classification (LAC) Dataset Construction
A.1 Preprocessing
@@
For the base dataset, we utilize the “evaluation” split of WikiBio Lebret et al. (2016), a personal biography dataset with information extracted from Wikipedia. We preprocess the dataset by splitting each personal biography into sentences. To ensure that each sentence is informative and depicts the owner of the biography, we remove the first two sentences and the last sentence, which usually provide the birth date and the current status of the owners without describing their characteristics.
A.2 ChatGPT Generation
@@
For each of the pre-processed sentences in personal biographies, we prompt the gpt-3.5-turbo-1106 version of ChatGPT with one-shot example Wang et al. (2020) to paraphrase it into an agentic version and a communal version. Specific prompt used in the dataset generation process is provided in Table 7. This guarantees the balance of the constructed dataset and prevents the propagation of pre-existing biases in the classifier training process.
@@ A.3 Human Re-Annotation
A.3.1 Post-processing
After the completion of human annotation on the language classification dataset, we conduct post-processing of the data by removing invalid data entries and aligning annotator agreements. We first remove all entries that are marked as “na” by either human annotator. Then, since the sentences are obtained by prompting ChatGPT to generate agentic or communal paraphrases, we treat the output categories as ChatGPT’s labeling of the data and align these labels with that of human annotators. For most cases where a majority vote exists, we utilize the label with the most votes as the gold label in the final dataset. For very few cases where both human annotators provide a distinct and different label from ChatGPT’s labeling, we invite a third human annotator to determine the final label in the dataset.
Split # Entries # Agentic # Communal
Train $2,979$ $1,523$ $1,456$
Valid $372$ $192$ $180$
Test $373$ $199$ $174$
Table 9: Dataset Statistics
Appendix B Language Agency Classifiers
B.1 Model Choices
We experiment with BERT and RoBERTa to build discriminative classifiers for language agency. For generative classifier, we experiment with the Reinforcement Learning with Human Feedback (RLHF)-tuned Llama2 for dialogue use cases Touvron et al. (2023). Below, we provide details on training and inferencing the models. For BERT and RoBERTa, we conduct full fine-tuning. For Llama2, we test with zero-shot prompting, one-shot prompting, and LoRA fine-tuning.
@@ B.1 Model Choices
Generative Model For the Llama2 generative model, we experiment with $4$ different settings: zero-shot prompting without definition, zero-shot prompting with definition, one-shot prompting with definition and an example, and parameter-efficient fine-tuning with LoRA Hu et al. (2021). For reproducibility, we provide the full prompts used to probe Llama2 in zero-shot and few-shot settings in Table 10. For LoRA fine-tuning, we use a learning rate of $5e^{-5}$ to train for $5$ epochs. During inference, we follow the default generation configuration to set top-p to $1.0$, tok-k to $50$, and temperature to $1.0$.
Model Size Training Accuracy F1
@@
String Matching N/A N/A $46.65$ $31.81$ $46.65$ $29.68$
Sentiment 66M N/A $52.28$ $41.35$ $52.28$ $43.05$
Llama2 7B + Base $82.56$ $49.46$ $50.38$ $50.03$
@@
Bert 109M + Fine-Tune 91.69 91.69 91.63 91.68
RoBERTa 125M + Fine-Tune $91.33$ $91.33$ $91.29$ $91.33$
Table 11: Classifier Statistics
B.2 Model Performance
Result of model performances on the proposed LAC dataset’s test set is reported in Table 11. To guarantee validity, all experiment results are averaged on random seeds 00, $1$, and $2$. For further experiments, we choose to use BERT model as the classifier since it achieves the highest test accuracy.
Appendix C Evaluation Dataset Details
In this study, we utilized $6$ datasets of human-written and LLM-synthesized texts. We briefly summarized the details and construction types of each dataset in Table  12, and provide additional information on data preprocessing and dataset synthesis below.
Dataset Type Note
Biography(Human) Human Bias in Bios
Biography(LLM) LLM Synthesized
Professor Review(Human) Human RateMyProfessor
Professor Review(LLM) LLM Synthesized
Reference Letter(Wan et al.) LLM  Wan et al. (2023)
Reference Letter(LLM) LLM Synthesized
Table 12: Details of 6 evaluation datasets.
C.1 Datasets in Previous Studies
@@
Bias in Bios For Bias in Bios De-Arteaga et al. (2019), we randomly sample $120$ biographies for each gender for each of the professions. Table 13 shows the full list of professions in the pre-processed dataset.
@@
RateMyProfessor We preprocess the RateMyProfessor dataset by removing data for departments where only less than $10$ reviews are available for male or female professors. Table 14 shows a full list of departments and the number of reviews for male and female professors under each department in the pre-processed dataset.
C.2 Generating New Datasets with ChatGPT
In addition to using previously proposed datasets, we also synthesize $1$ biography dataset, $1$ professor review dataset, and $1$ reference letter dataset using ChatGPT. Specifically, we adopt the Context-Less Generation approach proposed by  Wan et al. (2023) to prompt the model using descriptor-based prompt templates, with gender, race, and additional personal information such as name, occupation, age, and department. To identify natural names for males and females in different racial groups, we prompted ChatGPT to generate $10$ popular male names and $10$ popular female names in each of the $4$ race categories: “White”, “Black”, “Hispanic”, and “Asian”. Table 15 shows the full list of investigated racial groups and corresponding popular names for each gender. Table 16 presents lists of descriptors including occupations, departments, and ages. Table 17 demonstrates the prompt templates used to generate synthetic biographies, professor reviews, and reference letter datasets.
Popular Names in Racial Groups
Female Names "Emily", "Ashley", "Jessica", "Sarah", "Elizabeth"
Black Male Names "Jamal", "Malik", "Tyrone", "Xavier", "Rashad"
Female Names "Jasmine", "Aaliyah", "Keisha", "Ebony", "Nia"
Hispanic Male Names "Juan", "Alejandro", "Carlos", "José", "Diego"
Female Names "María", "Ana", "Sofia", "Gabriela", "Carmen"
Asian Male Names "Wei", "Hiroshi", "Minh", "Raj", "Jae-Hyun"
Female Names "Mei", "Aiko", "Linh", "Priya", "Ji-Yoon"
Table 15: Racial groups and popular male and female names for evaluation dataset synthesis.
Appendix D Additional Experiment Results

### v3

@@ 5 Conclusion
In this work, we propose the Language Agency Bias Evaluation (LABE) framework to systematically and comprehensively measure gender, racial, and intersectional biases in language agency across a wide scope of text generation tasks. To build better agency evaluation tools, we also contribute the Language Agency Classification (LAC) dataset for training accurate language agency classifiers. Through experimenting on 3 LLMs, we found that: (1) LLM-generated texts often carry remarkably higher levels of bias than human-written language; (2) People who are at the intersection of gender and racial minority groups (e.g. Black females) are the most vulnerable to language agency biases; (3) Compared with ChatGPT and Mistral, Llama3’s outputs tend to show the greatest overall level of language agency bias; (4) Simple prompt-based mitigation methods might result in the amplification and overshooting of biases, worsening the fairness issue in LLMs. Our LABE benchmark addresses previous research gaps and provides valuable technical contributions, and our findings point towards the imminence of comprehensively examining and resolving language agency biases in LLMs, forewarning potential social risks of using LLM-generated texts without scrutiny.
Appendix A Prompt Construction for Text Generation Tasks in LABE
We adapt the Context-Less Generation (CLG) approach proposed by  Wan et al. [2023] to prompt LLMs on biography, professor review, and reference letter generation tasks using descriptor-based prompt templates. All prompts contain gender and race descriptors to include demographic information. Additional information such as name, occupation, age, and department are also injected in prompts for different tasks. Table 5 demonstrates prompt templates used in LABE to obtain LLM outputs on the 3 text generation tasks about people with different demographic identities, for comprehensive bias analysis. Table 7 presents lists of descriptors including gender, race, occupations, departments, and ages. To identify natural names for males and females in different racial groups, we prompted ChatGPT to generate $5$ popular male names and $5$ popular female names in each of the $4$ race categories: “White”, “Black”, “Hispanic”, and “Asian”. Table 6 shows the full list of investigated racial groups and corresponding popular names for males and females within each group.
Text Generation Type Prompt # Entries
Biography "Generate a personal biography for {name}, a {age} year old {race} {gender} {occupation}." 2,400
Professor Review "Generate a review for {name}, a {race} {gender} professor in the {department name}." 600
Reference Letter "Generate a detailed reference letter for {name}, a {age} year old {race} {gender} {occupation}." 2,400
Table 5: Prompt templates for the 3 text generation tasks in LABE.
Race Gender Popular Names
Female Names "Emily", "Ashley", "Jessica", "Sarah", "Elizabeth"
Black Male Names "Jamal", "Malik", "Tyrone", "Xavier", "Rashad"
Female Names "Jasmine", "Aaliyah", "Keisha", "Ebony", "Nia"
Hispanic Male Names "Juan", "Alejandro", "Carlos", "José", "Diego"
Female Names "María", "Ana", "Sofia", "Gabriela", "Carmen"
Asian Male Names "Wei", "Hiroshi", "Minh", "Raj", "Jae-Hyun"
Female Names "Mei", "Aiko", "Linh", "Priya", "Ji-Yoon"
Table 6: Racial groups and popular male and female names as descriptors for constructing templated-based text generation prompts in LABE.
Descriptor Type Descriptor Items
Gender "male", "female"
Race "White", "Black", "Hispanic", "Asian"
Names See Table 6.
Occupations "student", "entrepreneur", "actor", "artist", "chef", "comedian", "dancer", "model", "musician", "podcaster", "athlete", "writer"
Departments "Communication department", "Fine Arts department", "Chemistry department", "Mathematics department", "Biology department", "English department", "Computer Science department", "Sociology department", "Economics department", "Humanities department", "Science department", "Languages department", "Education department", "Accounting department", "Philosophy department"
Ages 20, 30, 40, 50, 60
Table 7: Descriptors for evaluation dataset synthesis.
Appendix B Language Agency Classification (LAC) Dataset Construction
B.1 Preprocessing
@@
For the base dataset, we utilize the “evaluation” split of WikiBio Lebret et al. [2016], a personal biography dataset with information extracted from Wikipedia. We preprocess the dataset by splitting each personal biography into sentences. To ensure that each sentence is informative and depicts the owner of the biography, we remove the first two sentences and the last sentence, which usually provide the birth date and the current status of the owners without describing their characteristics.
B.2 ChatGPT Generation
@@
For each of the pre-processed sentences in personal biographies, we prompt the gpt-3.5-turbo-1106 version of ChatGPT with one-shot example Wang et al. [2020] to paraphrase it into an agentic version and a communal version. Specific prompt used in the dataset generation process is provided in Table 8. This guarantees the balance of the constructed dataset and prevents the propagation of pre-existing biases in the classifier training process.
@@ B.3 Human Re-Annotation
B.4 Post-processing
After the completion of human annotation on the language classification dataset, we conduct post-processing of the data by removing invalid data entries and aligning annotator agreements. We first remove all entries that are marked as “na” by either human annotator. Then, since the sentences are obtained by prompting ChatGPT to generate agentic or communal paraphrases, we treat the output categories as ChatGPT’s labeling of the data and align these labels with that of human annotators. For most cases where a majority vote exists, we utilize majority voting to determine the gold label in the final dataset. For very few cases where both human annotators provide a distinct and different label from ChatGPT’s labeling, we invite a third expert annotator to determine the final label in the dataset.
Appendix C Building Language Agency Classifiers with LAC
We provide details of training and inferencing language agency classification tools below. Licensing information for each model involved is provided in Table 11.
C.1 Model Choices
We experiment with BERT and RoBERTa to build discriminative classifiers for language agency. For generative classifier, we experiment with the Reinforcement Learning with Human Feedback (RLHF)-tuned Llama2 for dialogue use cases Touvron et al. [2023]. Below, we provide details on training and inferencing the models. For BERT and RoBERTa, we conduct full fine-tuning. For Llama2, we test with zero-shot prompting, one-shot prompting, and LoRA fine-tuning.
@@ C.1 Model Choices
Generative Model For the Llama2 generative model, we experiment with $4$ different settings: zero-shot prompting without definition, zero-shot prompting with definition, one-shot prompting with definition and an example, and parameter-efficient fine-tuning with LoRA Hu et al. [2021]. For reproducibility, we provide the full prompts used to probe Llama2 in zero-shot and few-shot settings in Table 10. For LoRA fine-tuning, we use a learning rate of $5e^{-5}$ to train for $5$ epochs. During inference, we follow the default generation configuration to set top-p to $1.0$, tok-k to $50$, and temperature to $1.0$.
Model Size License Training Accuracy F1
@@
String Matching N/A N/A N/A $46.65$ $31.81$ $46.65$ $29.68$
Sentiment 66M Apache 2.0 License N/A $52.28$ $41.35$ $52.28$ $43.05$
[Wan et al., 2023] 109M MIT License + Fine-Tune $66.49$ $66.49$ $64.22$ $64.82$
Llama2 7B LLAMA 2 Community License + Base $82.56$ $49.46$ $50.38$ $50.03$
@@
Bert 109M Apache 2.0 License + Fine-Tune 91.69 91.69 91.63 91.68
RoBERTa 125M MIT License + Fine-Tune $91.33$ $91.33$ $91.29$ $91.33$
Table 11: Performance details of different language agency classification methods. Licensing information specified for all models involved.
C.2 Model Performance
We report the performances of baseline methods to classify language agency, as well as our trained classifiers on the LAC dataset. For baseline methods, we experimented on string matching, sentiment classification, and the agency classifier proposed in Wan et al. [2023]’s work. For string matching, we utilized Stahl et al. [2022]’s released lists of agentic and communal words with no licensing information. For sentiment classification, we utilized the sentiment classification pipeline in the transformers library with the off-the-shelf “distilbert/distilbert-base-uncased-finetuned-sst-2-english” [^3] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english model.
Result of model performances on the proposed LAC dataset’s test set is reported in Table 11. Based on performance results, we choose to use BERT model as the classifier for further experiments since it achieves the highest test accuracy.
Appendix D Human-Written Datasets Details
In this study, we utilized $3$ datasets of human-written texts. We provide additional information on data preprocessing below.
Bias in Bios   The Bias in Bios De-Arteaga et al. [2019] dataset is released under MIT license. For preprocessing this dataset, we randomly sample $120$ biographies for each gender for each of the professions. Table 12 shows the full list of professions in the pre-processed dataset.
@@ Appendix D Human-Written Datasets Details
RateMyProfessor   The RateMyProfessor has no displayed licensing information and is publicly available on GitHub. We preprocess the RateMyProfessor dataset by removing data for departments where only less than $10$ reviews are available for male or female professors. Table 13 shows a full list of departments and the number of reviews for male and female professors under each department in the pre-processed dataset.
@@
Reference Letter [Wan et al., 2023] We directly utilize the reference letter dataset generated under the CBG setting in  Wan et al. [2023]’s work, which is released under the MIT License.
Appendix E Additional Experiment Results